OpenStudy (anonymous):

integrate 9dv/81 - v^4

OpenStudy (anonymous):

\[\int\limits_{}^{}9dv/81 - v^4\]

OpenStudy (anonymous):

is it integral 9/81 -v^4 dv?

OpenStudy (anonymous):

1/9-v^4 dv 1/9v-(v^5)/5+C

OpenStudy (anonymous):

\[\int\limits_{}^{}(9dv)/(81 - v^4)\]

OpenStudy (anonymous):

Oh !

OpenStudy (anonymous):

it uses tanx

OpenStudy (anonymous):

it is pretty lengthy

OpenStudy (anonymous):

my solutions manual shows \[1/12 \ln|(3+v)/(3-v)| + (1/6) \tan^{-1} (v/3) + C\]

OpenStudy (amistre64):

81-v^4 = (9+v^2)(9-v^2) = (9+v^2)(3+v)(3-v)

OpenStudy (amistre64):

id do partial decomps

OpenStudy (amistre64):

\[\frac{9}{(81-v^4)}=\frac{Ax+B}{(9+v^2)}+\frac{C}{(3-v)}\frac{D}{(3+v)}\]

OpenStudy (amistre64):

Av+B that is

OpenStudy (anonymous):

I KNEW IT WAS PARTIAL FRACTIONS!!!!!!

OpenStudy (amistre64):

lol .... they do help

OpenStudy (anonymous):

yea i was stuck on how i go about breaking it down to irreducible factors though

OpenStudy (anonymous):

but i knew it was somewhere with partial fractions thanks

OpenStudy (amistre64):

\[\frac{9}{(81-v^4)}=\frac{Av+B}{(9+v^2)}+\frac{C}{(3-v)}+\frac{D}{(3+v)}\] \[\small 9=Av+B(3-v)(3+v)+C(3+v)(9+v^2)+D(3-v)(9+v^2)\]

OpenStudy (anonymous):

for future reference, how do i place numbers over other numbers in the equations tab when i ask a question? Instead of using the / sign.